OpenStudy (anonymous):

what is the equations of the line passing through the points (-2,3) and (1,4)

OpenStudy (sidsiddhartha):

equation of a line passing through (x1,y1) and (x2,y2) is \[\frac{ x-x1 }{ y-y1 }=\frac{ x1-x2 }{ y1-y2 }\] here x1=-2,y1=3 ,x2=1,y2=4